REU 2008: SCHEDULE FOR WEEKS TWO AND THREE

Date  YSP  APP/SES  Abert  Babai  Farb  May  Dicksons*
Mon, June 30 9:30–2:30 9:30–12:00 2:45–4:15  4:30–5:30  TBA
Tue, July 1  9:30–2:30 9:30–12:00  4:30–5:30 2:45–4:15  TBA
Wed, July 2  9:30–2:30 9:30–12:00 2:45–4:15  4:30–5:30  TBA
Thu, July 3  9:30–2:30 9:30–12:00  4:30–5:30 2:45–4:15  TBA
Date  YSP  APP/SES  Abert  Farb  May  Dicksons*
Mon, July 7 9:30–2:30 9:30–12:00 2:45–4:15  4:30–5:30  TBA
Tue, July 8 9:30–2:30 9:30–12:00  2:45–4:15 4:30–5:30  TBA
Wed, July 9 9:30–2:30 9:30–12:00 2:45–4:15  4:30–5:30  TBA
Thu, July 10 9:30–2:30 9:30–12:00  2:45–4:15 4:30–5:30  TBA
Fri, July 11 9:30–2:00 9:30–12:00 2:15–3:45  4:00–5:00  TBA

* Angeltveik, Basak, Crans, McReynolds

Here is the schedule in summary form:

SECOND WEEK:

YSP: 9:30–2:30 M–Th Abert/Babai (Apprentice/SESAME people welcome): 9:30–12:00 M–Th, R251 Abert (Problems): 2:45–4:15 MW, R251 Babai (Discrete mathematics): 4:30–5:30 TTh, E133 Farb (Topics in geometry): 2:45–4:15 TTh, R251 May (Finite spaces): 4:30–5:30 MW, E133 Postdoc sessions in geometry and topology: TBA

THIRD WEEK:

YSP: 9:30–2:30 M–Th, 9:30–2:00 F Abert (Apprentice/SESAME people welcome): 9:30–12:00 M–F, R251 Abert (Problems): 2:45–4:15 MW, 2:15–3:45 F, E133 Farb (Topics in Geometry): 2:45–4:15 TTh, R251 May (Finite Spaces): 4:30–5:30 M–TH, 4:00–5:00 F, E133 Postdoc sessions in geometry and topology: TBA